40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10659     Entry time: Fri Oct 31 19:59:26 2014
Author: Koji 
Type: Update 
Category: General 
Subject: Some locking work / PRMI analysis 

Preparations

- According to Diego's report, the MC WFS gains were too high. We'll fix this later by tweaking the servo shapes.
But for now, all of the WFS gains were reduced by 40%.
i.e. WFS(1|2)(PIT|YAW) gains from 5 to 3, MC2TRANS(PIT|YAW) gains from 50 to 30.

- Aligned IMC carefully and ran the offset nulling script. MC REFL became 0.435~0.445 and MC TRANS was ~16600.

- Locked the arms and ran ASS.


PRMI

- Started locking PRMI. I just used REFL33I&Q as suggested by the configure script. The PRMI locking was not so robust.
Particularly, the third violin mode of PRM and BS seemed to get excited and dominated the signals.
I modified Vio3 filter in the violin filter for BS and PRM to include zero at 1921Hz where the growing peak was seen.

- We probably want to start from the 1f signals for DRMI lock acquisition. So I wanted to check how REFL11s are.
Measured the demod phase and relative gain between 33I and 11I. (By the way, REFL11I whitening gain was lowered to 0dB).
REFL11I had about x10 gain and the same phase compared to REFL33I. The demod phase for REFL11 was +21deg.
Also checked REFL55 phase and gain. 55Q has almost the same gain as 33Q. And the adjusted phase was 25deg.
These were just rough adjustment of the demod phases.

- Then the servo configuration was transtioned to Configuration 1 (below), and then Configuration 2.

- This configuration was very stable and the PRMI stayed locked about ~1 hour. During this long lock, I could measure 
PSDs, sensing matrix, and etc. Also I could play with the PRM ASC. I wasn't sure if the POP is actually stabilized or not.
(I have no data)

- I noticed that something was ringinging up at 1883Hz. Another 3rd order viloin mode???

- The lock was lost due to too strong injection. But also it reacquired without touching.

- Precise demod phase adjustment has been done by elliminating PRCL from the Q signals.

REFL11 16.75
REFL33 133.0
REFL55 31.0
REFL165 -142 
AS55 -53

- Configiration1 (REFL11I&REFL55Q)

REFL11: WTN 0dB PHASE 21deg, REFL11I x0.1 -> PRCL
REFL33: WTN 30dB PHASE 145deg
REFL55: WTN 21dB PHASE 25deg, REFL55Q x1 -> MICH

PRCL: GAIN -0.04 FM4/5 ON, Triggered FM 2/3/6/9, Servo trigger: POP22I 50up 10down, No Normaization.
MICH: GAIN 10 FM4/5 ON, Triggered FM 2/3/6/9, Servo trigger: POP22I 50up 10down, No Normaization.

PRCL -> PRM +1
MICH -> PRM -0.2625, BS +0.50 BS

- Configuration 2 (REFL11I&Q)

Same as above except:
REFL11Q x-0.1 -> MICH


Calibration

Let's use these entries 

PRM: http://nodus.ligo.caltech.edu:8080/40m/8255
PRM:  (19.6 +/- 0.3) x 10^{-9} (Hz/f)^2 m/counts

BS/ITMs http://nodus.ligo.caltech.edu:8080/40m/8242
BS     = (20.7 +/- 0.1)    x 10 -9 / f2
ITMX = (4.70 +/- 0.02)  x 10 -9/ f2
ITMY = (4.66 +/- 0.02) x 10 -9/ f2

- PRCL Calibration

Lockin oscillator module 675.13Hz 100 -> +1 PRM

Measurement bandwidth 0.1Hz -> Signal power BW 0.471232 (FLATTOP window)

C1:SUS-PRM_LSC_IN1: 118.99 cnt/rtHz => 5.12pm/rtHz
REFL11I: 17.84  cnt/rtHz => 3.49e12 cnt/m
REFL33I:  2.28  cnt/rtHz => 4.46e11 cnt/m
REFL55I:  0.158 cnt/rtHz => 3.09e10 cnt/m
REFL165I: 1.63  cnt/rtHz => 3.19e11 cnt/m


- MICH Calibration

Lockin oscillator module 675.13Hz 100 -> -1 ITMX +1 ITMY

Measurement bandwidth 0.1Hz -> Signal power BW 0.471232 (FLATTOP window)

C1:SUS-ITMX_LSC_IN1: 121.79 cnt/rtHz => 1.26pm/rtHz
C1:SUS-ITMY_LSC_IN1: 121.79 cnt/rtHz => 1.25pm/rtHz
REFL11Q:  0.0329   cnt/rtHz => 1.32e10 cnt/m (PRCL/MICH ratio 265)
REFL33Q:  0.00773  cnt/rtHz => 3.09e9  cnt/m (144)
REFL55Q:  0.001645 cnt/rtHz => 6.58e8  cnt/m (47)
REFL165Q: 0.00374  cnt/rtHz => 1.50e9  cnt/m (213) !?
AS55Q:    0.0696   cnt/rtHz => 2.78e10 cnt/m

Openloop TF measurements
Servo filter TF measuremnts

The UGFs were ~250Hz for PRCL and ~120Hz for MICH, respectively.
The OLTF was modelled by the servo and violin filters TF from foton, estimated TF of the AA/AI filters, and the constant time delay.

Displacement spectra measurement

SELF NOTE: DON'T FORGET TO TURN ON the whitening of the unused signals! (USE MC DOF or manual switch)

- PRCL

The PRCL displacement was measured with REFL I signals. In the attachment 3, the in-loop and free-run equivalent displacements are shown (red and blue).
Other out-of-loop sensors (33/55/165) were also plotted together.

FIrst of all, the uncompensated displacement noise level of PRCL is around 1e-7 m/rtHz. This is a good indication that the calibration was not crazy.

The sensing noise of REFL11 seems to be 1e-15~1e-16 m/rtHz at high frequency which is enough for now.
As expected, REFL11I has the best noise level among the REFLs. At low frequency, it seemed that the noise level is limited by something at 1e-12 m/rtHz.
Of course, we can't say this is just the sensing noise of the other REFLs or the noise of the REFL11I. But this noise level is enough small for the locking of
the low finesse (F<100) PRCL cavity.

Remembering we had no trouble locking PRCL with REFL33/55/165, this plot indicates that the PRCL was suppressed too much below 2Hz.
And we want more supression between 5Hz to 30Hz. We have resonant gains in ther PRCL servo but not sure how effective they were.
If we consider the contamination of PRCL in MICH, we should try to optimize the PRCL servo.

- MICH

The MICH displacement was similary calibrated to PRCL. The signal sources were the REFL Qs and AS55Q.
In the attachment 4, the in-loop and free-run equivalent displacements are shown (red and blue).
Other out-of-loop sensors were also plotted together.

The problem here is that the out-of-loop levels (REFL33/55/165 and AS55) show almost the same levels
and thus it is likely that the actual (out-of-loop) stability of MICH is this kind of level. If we believe it, we only have
~1/100 supression between 1-10Hz and ~1/10Hz below 0.5Hz.
The strong servo control does nothing to stablize
MICH. From the out-of-loop noise level of MICH, this comes for the contamination from leakage PRCL.
We really need to improve the signal quality of MICH.

The MICH servo filter has quite complicated shape, but is not necessary according to the estimated free-runing MICH.

The MICH free-running motion is quieter than the PRCL one between 1Hz to 30Hz. The reasonable explanation is
that it comes from poor vibration isolation of the tip-tilts. It means that SRCL also has the similar noise level to PRCL.

Attachment 1: PRMIsb_PRCL_OLTF.pdf  1.372 MB  Uploaded Mon Nov 3 13:44:42 2014  | Hide | Hide all
PRMIsb_PRCL_OLTF.pdf
Attachment 2: PRMIsb_MICH_OLTF.pdf  1.380 MB  Uploaded Mon Nov 3 13:45:02 2014  | Hide | Hide all
PRMIsb_MICH_OLTF.pdf
Attachment 3: PRMIsb_PRCL_SPE.pdf  1.931 MB  Uploaded Mon Nov 3 13:45:20 2014  | Hide | Hide all
PRMIsb_PRCL_SPE.pdf
Attachment 4: PRMIsb_MICH_SPE.pdf  1.956 MB  Uploaded Mon Nov 3 13:45:41 2014  | Hide | Hide all
PRMIsb_MICH_SPE.pdf
ELOG V3.1.3-